1, you are offered to receive $200, $250, $300,$350 and $400 at the end of each year the first payment of $200 is giving to you two years from now. if your money can earn 5% annual rate. how much you should pay for this deal?

2, your father will retire in 10 years. if he needs $25,000 per years for 10 years after retirement. how much he have to save at the end of each year in the next 10 years to support your after retirement needs if market interest rate for saving is on average is 5%annually for next 20 years.

3, you friend is going to purchase a car. the cost of a car is $22,000. she put $5,000 as down payment and borrow the rest from dealer. The dealer finance is a 6% APR, with monthly payment for 48 months. what is your friend's monthly payment?
Answer #1

­SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASEW i v a 1x 3 € ENG 03:02 21-03-2020 . O x M383 X ✓ fr L M N O P Q R S T A 366 367 368 369 YEAR END CASH FLOWS PV FACTOR AT 5%

= W @ vera 1x 3 ENG 1 03:04 21-03-2020 AE253 - X fix AD AE AF AG AH AI AJ AK A 232 233 234 RETIREMENT BA II PLUS SET C/Y=P/Y

@ vera 1x 3 ENG 03:06 W F 21-03-2020 ... o X A263 - fix X B A C F G H I 251 252 253 set C/Y=P/Y =12 254 255 N = 48 256 257 -1